# expected utility

### v2

@@ Appendix A Proofs
    Since $\lim_{\lambda\searrow C}h(\lambda)=0$ and $C>-2\gamma\alpha k_{0}$, it holds that (i) $\xi_{3}^{*}\to+\infty$ for $\lambda\searrow C$, i.e., $y=h(\lambda)\xrightarrow{\lambda\searrow C}0$. Indeed, it holds: If $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}>k_{2}$, we get that $C>2\gamma\alpha(k_{2}-k_{1}-k_{0})$ and hence $\hat{\xi}\xrightarrow{\lambda\searrow C}+\infty$ since $y=h(\lambda)\xrightarrow{\lambda\searrow C}0$. Now, (A.16) implies that $\xi_{3}^{*}\xrightarrow{\lambda\searrow C}+\infty$. If $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}\leq k_{2}$, it holds that $C=2\gamma\alpha\left(e^{\int_{0}^{T}r_{s}\mathrm{d}s}-k_{1}-k_{0}\right)$ since $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}>k_{1}$ by assumption and we get that:
@@
    Now, we see that $\lim_{\lambda\searrow C}D>0$ if and only if $\left(x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}-k_{1}\right)^{2}>0$ which is true by assumption. Thus, $\xi_{3}^{*}\xrightarrow{\lambda\searrow C}+\infty$ since $h(\lambda)\xrightarrow{\lambda\searrow C}0$, i.e., (i) is proven.
@@
    Summarizing, we get in all subcases of Case 1, that $\Phi(d_{0}(\xi_{3}^{*},0))\xrightarrow{\lambda\searrow C}1$ since $\lim_{x\to\infty}\Phi\left(d_{0}\left(x,0\right)\right)=1$ due to (i) and $-\Phi(d_{0}(\alpha\hat{\xi}^{\lambda_{n}},0))+\Phi(d_{0}(\xi_{2}^{*,\lambda_{n}},0))-\Phi(d_{0}(\tilde{\alpha}\hat{\xi}^{\lambda_{n}},0))+\Phi(d_{0}(\xi_{1}^{*,\lambda_{n}},0))\xrightarrow{n\to\infty}0$ due to either (ii), (iii), (iv), (v), or (vi) depending on $C$. Hence, it holds that $\Phi(d_{0}(\xi_{3}^{*,\lambda_{n}},0))-\Phi(d_{0}(\alpha\hat{\xi}^{\lambda_{n}},0))+\Phi(d_{0}(\xi_{2}^{*,\lambda_{n}},0))-\Phi(d_{0}(\tilde{\alpha}\hat{\xi}^{\lambda_{n}},0))+\Phi(d_{0}(\xi_{1}^{*,\lambda_{n}},0))\xrightarrow{n\to\infty}1$ since $\lim_{x\to\infty}\Phi\left(d_{0}\left(x,0\right)\right)=1$ and $\lambda_{n}\xrightarrow{n\to\infty}C$ from above. In particular, we get that $\lambda(-1+\Phi(d_{0}(\xi_{3}^{*,\lambda_{n}},0))-\Phi(d_{0}(\alpha\hat{\xi}^{\lambda_{n}},0))+\Phi(d_{0}(\xi_{1}^{*,\lambda_{n}},0)))+2\gamma\alpha(k_{2}-k_{1}-k_{0})(\Phi(d_{0}(\xi_{2}^{*,\lambda_{n}},0))-\Phi(d_{0}(\tilde{\alpha}\hat{\xi}^{\lambda_{n}},0)))\xrightarrow{n\to\infty}0$ since $C=2\gamma\alpha(k_{2}-k_{1}-k_{0})$ and $\lambda_{n}\xrightarrow{n\to\infty}C$ from above. Thus, the claim follows.
@@
    It holds that $\bar{\xi}^{\lambda_{n}}\xrightarrow{n\to\infty}0$ using $y_{n}=h(\lambda_{n})$ and the assumption. Therefore, we also get that $\hat{\xi}^{\lambda_{n}},\xi_{1}^{*,\lambda_{n}},\xi_{2}^{*,\lambda_{n}},\xi_{3}^{*,\lambda_{n}}\xrightarrow{n\to\infty}0$ due to (A.16). Thus, the claim follows using $\lim_{x\to 0}\Phi(d_{i}(x,0))=0$ for $i\in\{0,1\}$.
@@
    To show this, it is sufficient to show that (i) $\lambda\mapsto f_{1}(\cdot,\lambda)$ is strictly increasing, (ii) $y\mapsto f_{1}(y,\cdot)$ is non-increasing and (iii) $f_{1}$ is jointly continuous on $\mathbb{R}_{>0}\times(C,\infty)$. Indeed, let $\lambda,\lambda_{n}>C$ with $\lambda_{n}\xrightarrow{n\to\infty}\lambda$. Due to the existence of the zero root (see the main part of the proof) and the uniqueness of the zero root (since $\lambda\mapsto f_{1}(\cdot,\lambda)$ is strictly increasing), there exist then unique $y_{\lambda}^{*},y_{\lambda_{n}}^{*}\in(0,\infty)$ such that $f_{1}(y_{\lambda}^{*},\lambda)=0=f_{1}(y_{\lambda_{n}}^{*},\lambda_{n})$. Then, we have to show that $y_{\lambda_{n}}^{*}\xrightarrow{n\to\infty}y_{\lambda}^{*}$. Due to $y\mapsto f_{1}(y,\cdot)$ being non-increasing and $\lambda_{n}\xrightarrow{n\to\infty}\lambda$, it holds that $y_{\lambda_{n}}^{*}\in[y_{\max_{k\in\mathbb{N}}\{\lambda,\lambda_{k}\}}^{*},y_{\min_{k\in\mathbb{N}}\{\lambda,\lambda_{k}\}}^{*}]$ (maximum and minimum exist due to the convergence of $\lambda_{n}$ to $\lambda$). Then, there exist a subsequence $y_{\lambda_{n_{l}}}^{*}$ and a $\tilde{y}$ such that $y_{\lambda_{n_{l}}}^{*}\xrightarrow{l\to\infty}\tilde{y}$. Moreover, we know that $f_{1}(y_{\lambda}^{*},\lambda)=0=f_{1}(y_{\lambda_{n_{l}}}^{*},\lambda_{n_{l}})\xrightarrow{l\to\infty}f_{1}(\tilde{y},\lambda)$ due to the joint continuity of $f_{1}$. Now, the uniqueness of the zero root implies that $\tilde{y}=y_{\lambda}^{*}$. Hence, all subsequences converge to $y_{\lambda}^{*}$ and thus also the sequence itself. Therefore, (A.12) would follow provided we can show (i), (ii), and (iii) from the beginning:
@@
    Indeed, if for all $\omega\in\Omega$, we can show that $\hat{X}_{T}^{\lambda_{1}}(\xi_{T}(\omega))\geq\hat{X}_{T}^{\lambda_{2}}(\xi_{T}(\omega))$ with a strict inequality for a set with positive probability for all $\lambda_{1}>\lambda_{2}\,(>C)$, then also $f_{1}(\cdot,\lambda_{1})>f_{1}(\cdot,\lambda_{2})$ since $f_{1}=\mathbb{E}[\xi_{T}\hat{X}_{T}]-x_{0}$ and $\xi_{T}>0$. Therefore, let $\lambda_{1}>\lambda_{2}\,(>C)$: Due to (A.17) and Proposition 3.4, we conclude that for all $\lambda>C$ (and hence for $\lambda_{1}$ and $\lambda_{2}$) $\hat{X}_{T}\not\equiv 0$, i.e., $\hat{X}_{T}(\xi_{T})>0$ for $\xi_{T}$ small enough. The formula of $\hat{X}_{T}$ (see Theorem 3.2) implies that for fixed $y$ the slope remains unchanged in each interval $(0,\xi_{1}^{*}]$, $(\tilde{\alpha}\hat{\xi},\xi_{2}^{*}]$, resp. $(\alpha\hat{\xi},\xi_{3}^{*}]$ when changing $\lambda$, but the interval boundaries change. Note that the slope (as a function of $\xi_{T}$) is strictly negative in $(0,\xi_{1}^{*}]$ and $(\alpha\hat{\xi},\xi_{3}^{*}]$, and constant otherwise. Therefore, due to $\hat{X}_{T}$ being non-increasing and a non-increasing function getting larger when being shifted to the right, it is sufficient to show that all interval boundaries do not decrease when $\lambda$ increases and at least one boundary value strictly increases when $\lambda$ increases:
@@
    Third, we prove property (ii), i.e., that $f_{1}$ is non-increasing in $y$. Let $\hat{X}_{T}(y)$ be a function of $y$ as in Theorem 3.2. It follows directly from the definitions of $\hat{\xi}$, $\xi_{1}^{*}$, $\xi_{2}^{*}$, and $\xi_{3}^{*}$ that they are non-increasing in $y$ for fixed $\lambda$. Now, let $y_{1}<y_{2}$ and $\xi_{T}>0$ arbitrary. Then, the claim follows if $\hat{X}_{T}(y_{1},\xi_{T})\geq\hat{X}_{T}(y_{2},\xi_{T})$. If $\xi_{T}$ is in the same interval for both values $y_{1}$ and $y_{2}$, the claim is obvious due to (A.7). Since $\hat{\xi}$, $\xi_{1}^{*}$, $\xi_{2}^{*}$, and $\xi_{3}^{*}$ are non-increasing in $y$ and $\hat{X}_{T}$ is non-increasing in $\xi_{T}$ (check Remark 3.3(c)), the claim also follows if $\xi_{T}$ is in different intervals for $y_{1}$ and $y_{2}$. ∎
@@
    Now, if $\limsup_{\varepsilon\to 0}y_{\lambda_{\varepsilon}}^{*}>0$, it follows that $\liminf_{\varepsilon\to 0}f_{1}(y_{\lambda_{\varepsilon}}^{*},\lambda_{\varepsilon})<0$ (since $\Phi\left(d_{1}\left(\xi_{3}^{*},0\right)\right)<1$ for all $y>0$) which is a contradiction to (A.8). Thus, the claim follows, i.e., $\lim_{\lambda\to C}y_{\lambda}^{*}=0$, when taking the limit of $\varepsilon\to 0$ (and hence $\lambda\to C$) on both sides, since $y_{\lambda}^{*}\geq 0$. Note that for every $y$, the existence of a solution $\lambda_{y}$ of (Proof of (A.13).) was already ensured in the main part of the proof.
@@
    $\leq$ $-x_{0}+x_{0}+\varepsilon-\displaystyle\frac{y}{2\gamma\tilde{\alpha}^{2}}e^{\int_{0}^{T}-2r_{s}+\left\lVert\kappa_{s}\right\rVert^{2}\mathrm{d}s},$
    where we used that $k_{2}e^{-\int_{0}^{T}r_{s}\mathrm{d}s}\leq x_{0}$, $\tilde{\alpha}+\alpha_{2}=\alpha$, and $\frac{\tilde{\alpha}}{\alpha}\leq 1$ in the first inequality. Moreover, we used (A.16) and $\Phi\left(d_{1}\left(\xi_{3}^{*},0\right)\right)\leq 1$ for all $y>0$ in the second inequality. Hence, the claim follows, i.e., $\lim_{\varepsilon\to 0}y_{\lambda_{\varepsilon}}^{*}=0$ by the same argument as in Case 1. Note that the existence of a solution was already ensured in the main part of the proof. ∎
@@
    We prove this by contradiction. Therefore, we assume that there exists an $L>0$ such that $\limsup_{\lambda\to\infty}y^{*}_{\lambda}\leq L$. Under this assumption, it holds that $\xi_{1}^{*}\xrightarrow{\lambda\to\infty}\infty$ since $\lambda-\sqrt{l(\lambda)}\xrightarrow{\lambda\to\infty}\infty$ for all affine functions $l$ and $y_{\lambda}^{*}$ being bounded. Then (A.16) implies that $\hat{\xi},\bar{\xi},\xi_{2}^{*},\xi_{3}^{*}\xrightarrow{\lambda\to\infty}\infty$ and we know that $\Phi(d_{1}(\xi_{3}^{*},0))\geq\Phi(d_{1}(\alpha\hat{\xi},0))$. However, then $\lim_{\lambda\to\infty}f_{1}(y,\lambda)=\infty$ for all $y\leq L$ which is a contradiction to $y^{*}_{\lambda}$ being the zero root of $f_{1}$. ∎

### v3

@@ Appendix A Proofs
    Since $\lim_{\lambda\searrow C}h(\lambda)=0$ and $C>-2\gamma\alpha k_{0}$, it holds that (i) $\xi_{3}^{*}\to+\infty$ for $\lambda\searrow C$, i.e., $y=h(\lambda)\xrightarrow{\lambda\searrow C}0$. Indeed, it holds: If $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}>k_{2}$, we get that $C>2\gamma\alpha(k_{2}-k_{1}-k_{0})$ and hence $\hat{\xi}\xrightarrow{\lambda\searrow C}+\infty$ since $y=h(\lambda)\xrightarrow{\lambda\searrow C}0$. Now, (A.17) implies that $\xi_{3}^{*}\xrightarrow{\lambda\searrow C}+\infty$. If $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}\leq k_{2}$, it holds that $C=2\gamma\alpha\left(x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}-k_{1}-k_{0}\right)$ since $x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}>k_{1}$ by assumption and we get that:
@@
    Now, we see that $\lim_{\lambda\searrow C}D>0$ if and only if $\left(x_{0}e^{\int_{0}^{T}r_{s}\mathrm{d}s}-k_{1}\right)^{2}>0$ which is true by assumption. Thus, $\xi_{3}^{*}\xrightarrow{\lambda\searrow C}+\infty$ since $y=h(\lambda)\xrightarrow{\lambda\searrow C}0$, i.e., (i) is proven.
@@
    Summarizing, we get in all subcases of Case 1 with $C=2\gamma\alpha(k_{2}-k_{1}-k_{0})$, that $\Phi(d_{0}(\xi_{3}^{*},0))\xrightarrow{\lambda\searrow C}1$ since $\lim_{x\to\infty}\Phi\left(d_{0}\left(x,0\right)\right)=1$ due to (i) and $-\Phi(d_{0}(\alpha\hat{\xi}^{\lambda_{n}},0))+\Phi(d_{0}(\xi_{2}^{*,\lambda_{n}},0))-\Phi(d_{0}(\tilde{\alpha}\hat{\xi}^{\lambda_{n}},0))+\Phi(d_{0}(\xi_{1}^{*,\lambda_{n}},0))\xrightarrow{n\to\infty}0$ due to either (ii), (iii), (iv), (v), or (vi) depending on $C$. Hence, it holds that $\Phi(d_{0}(\xi_{3}^{*,\lambda_{n}},0))-\Phi(d_{0}(\alpha\hat{\xi}^{\lambda_{n}},0))+\Phi(d_{0}(\xi_{2}^{*,\lambda_{n}},0))-\Phi(d_{0}(\tilde{\alpha}\hat{\xi}^{\lambda_{n}},0))+\Phi(d_{0}(\xi_{1}^{*,\lambda_{n}},0))\xrightarrow{n\to\infty}1$ since $\lim_{x\to\infty}\Phi\left(d_{0}\left(x,0\right)\right)=1$ and $\lambda_{n}\xrightarrow{n\to\infty}C$ from above. In particular, we get that $\lambda(-1+\Phi(d_{0}(\xi_{3}^{*,\lambda_{n}},0))-\Phi(d_{0}(\alpha\hat{\xi}^{\lambda_{n}},0))+\Phi(d_{0}(\xi_{1}^{*,\lambda_{n}},0)))+2\gamma\alpha(k_{2}-k_{1}-k_{0})(\Phi(d_{0}(\xi_{2}^{*,\lambda_{n}},0))-\Phi(d_{0}(\tilde{\alpha}\hat{\xi}^{\lambda_{n}},0)))\xrightarrow{n\to\infty}0$ since $C=2\gamma\alpha(k_{2}-k_{1}-k_{0})$ and $\lambda_{n}\xrightarrow{n\to\infty}C$ from above. Thus, the claim follows.
@@
    It holds that $\bar{\xi}^{\lambda_{n}}\xrightarrow{n\to\infty}0$ using $y_{n}=h(\lambda_{n})$ and the assumption. Therefore, we also get that $\hat{\xi}^{\lambda_{n}},\xi_{1}^{*,\lambda_{n}},\xi_{2}^{*,\lambda_{n}},\xi_{3}^{*,\lambda_{n}}\xrightarrow{n\to\infty}0$ due to (A.17). Thus, the claim follows using $\lim_{x\to 0}\Phi(d_{i}(x,0))=0$ for $i\in\{0,1\}$ and $h(\lambda_{n})\xrightarrow{n\to\infty}+\infty$.
@@
    To show this, it is sufficient to show that (i) $\lambda\mapsto f_{1}(\cdot,\lambda)$ is strictly increasing, (ii) $y\mapsto f_{1}(y,\cdot)$ is strictly decreasing and (iii) $f_{1}$ is jointly continuous on $\mathbb{R}_{>0}\times(C,\infty)$. Indeed, let $\lambda,\lambda_{n}>C$ with $\lambda_{n}\xrightarrow{n\to\infty}\lambda$. Due to the existence of the zero root (see the main part of the proof) and the uniqueness of the zero root (by (ii)), there exist then unique $y_{\lambda}^{*},y_{\lambda_{n}}^{*}\in(0,\infty)$ such that $f_{1}(y_{\lambda}^{*},\lambda)=0=f_{1}(y_{\lambda_{n}}^{*},\lambda_{n})$. Then, we have to show that $y_{\lambda_{n}}^{*}\xrightarrow{n\to\infty}y_{\lambda}^{*}$. Due to (i), (ii), and $\lambda_{n}\xrightarrow{n\to\infty}\lambda$, it holds that $y_{\lambda_{n}}^{*}\in[y_{\max_{k\in\mathbb{N}}\{\lambda,\lambda_{k}\}}^{*},y_{\min_{k\in\mathbb{N}}\{\lambda,\lambda_{k}\}}^{*}]$ (maximum and minimum exist due to the convergence of $\lambda_{n}$ to $\lambda$). Then, there exist a subsequence $y_{\lambda_{n_{l}}}^{*}$ and a $\tilde{y}$ such that $y_{\lambda_{n_{l}}}^{*}\xrightarrow{l\to\infty}\tilde{y}$. Moreover, we know that $f_{1}(y_{\lambda}^{*},\lambda)=0=f_{1}(y_{\lambda_{n_{l}}}^{*},\lambda_{n_{l}})\xrightarrow{l\to\infty}f_{1}(\tilde{y},\lambda)$ due to the joint continuity of $f_{1}$. Now, the uniqueness of the zero root implies that $\tilde{y}=y_{\lambda}^{*}$. Hence, all subsequences converge to $y_{\lambda}^{*}$ and thus also the sequence itself. Therefore, (A.13) would follow provided we can show (i), (ii), and (iii) from the beginning:
@@
    Indeed, if, for all $\omega\in\Omega$, we can show that $\hat{X}_{T}^{\lambda_{1}}(\xi_{T}(\omega))\geq\hat{X}_{T}^{\lambda_{2}}(\xi_{T}(\omega))$ with a strict inequality for a set with positive probability for all $\lambda_{1}>\lambda_{2}\,(>C)$, then also $f_{1}(\cdot,\lambda_{1})>f_{1}(\cdot,\lambda_{2})$ since $f_{1}=\mathbb{E}[\xi_{T}\hat{X}_{T}]-x_{0}$ and $\xi_{T}>0$. Therefore, let $\lambda_{1}>\lambda_{2}\,(>C)$: Due to (A.18) and Proposition 3.4, we conclude that for all $\lambda>C$ (and hence for $\lambda_{1}$ and $\lambda_{2}$) $\hat{X}_{T}\not\equiv 0$, i.e., $\hat{X}_{T}(\xi_{T})>0$ for $\xi_{T}$ small enough. The formula of $\hat{X}_{T}$ (see Theorem 3.2) implies that for fixed $y$ the slope remains unchanged in each interval $(0,\xi_{1}^{*}]$, $(\tilde{\alpha}\hat{\xi},\xi_{2}^{*}]$, resp. $(\alpha\hat{\xi},\xi_{3}^{*}]$ when changing $\lambda$, but the interval boundaries change. Note that the slope (as a function of $\xi_{T}$) is strictly negative in $(0,\xi_{1}^{*}]$ and $(\alpha\hat{\xi},\xi_{3}^{*}]$, and constant otherwise. Therefore, due to $\hat{X}_{T}$ being non-increasing and a non-increasing function getting larger when being shifted to the right, it is sufficient to show that all interval boundaries do not decrease when $\lambda$ increases and at least one boundary value strictly increases when $\lambda$ increases:
@@
    Third, we prove property (ii), i.e., that $f_{1}$ is strictly decreasing in $y$. Let $\hat{X}_{T}(y)$ be a function of $y$ as in Theorem 3.2. It follows directly from the definitions of $\hat{\xi}$, $\xi_{1}^{*}$, $\xi_{2}^{*}$, and $\xi_{3}^{*}$ that they are non-increasing in $y$ for fixed $\lambda$ since they are non-negative by definition and $\tfrac{1}{y}$ can be factored out. Now, let $y_{1}<y_{2}$ and $\xi_{T}>0$ arbitrary. Then, the claim follows if $\hat{X}_{T}(y_{1},\xi_{T})\geq\hat{X}_{T}(y_{2},\xi_{T})$ and $\hat{X}_{T}(y_{1},\xi_{T})>\hat{X}_{T}(y_{2},\xi_{T})$ on a set with positive probability. If $\xi_{T}$ is in the same interval for both values $y_{1}$ and $y_{2}$, $\geq$” is obvious due to (3.4). Moreover, we get ”$>$” if $\xi_{T}\in(0,\xi_{1}^{*}]\cup(\alpha\hat{\xi},\xi_{3}^{*}]$ (attained with positive probability). Since $\hat{\xi}$, $\xi_{1}^{*}$, $\xi_{2}^{*}$, and $\xi_{3}^{*}$ are non-increasing in $y$, $\hat{X}_{T}$ is non-increasing in $\xi_{T}$ (check Proposition 3.5), and for all $y\geq 0$ holds that $\hat{X}_{T}(y,a)>\hat{X}_{T}(y,b)>\hat{X}_{T}(y,c)>\hat{X}_{T}(y,d)$ for $a\in(0,\xi_{1}^{*})$, $b\in(\tilde{\alpha}\hat{\xi},\xi_{2}^{*}]$, $c\in(\alpha\hat{\xi},\xi_{3}^{*})$, and $d>\xi_{3}^{*}$, the claim also follows if $\xi_{T}$ is in different intervals for $y_{1}$ and $y_{2}$. ∎
@@
    Now, if $\limsup_{\varepsilon\to 0}y_{\lambda_{\varepsilon}}^{*}>0$, it follows that $\liminf_{\varepsilon\to 0}f_{1}(y_{\lambda_{\varepsilon}}^{*},\lambda_{\varepsilon})<0$ (since $\Phi\left(d_{1}\left(\xi_{3}^{*},0\right)\right)<1$ for all $y>0$) which is a contradiction to $0=\liminf_{\varepsilon\to 0}0=\liminf_{\varepsilon\to 0}f_{1}(y_{\lambda_{\varepsilon}}^{*},\lambda_{\varepsilon})$ by definition of $y_{\lambda_{\varepsilon}}^{*}$. Thus, the claim follows, i.e., $\lim_{\lambda\to C}y_{\lambda}^{*}=0$, when taking the limit of $\varepsilon\to 0$ (and hence $\lambda\to C$) on both sides, since $y_{\lambda}^{*}\geq 0$. Note that for every $y$, the existence of a solution $\lambda_{y}$ of (Proof of (A.14).) was already ensured in the main part of the proof.
@@
    $\leq$ $-x_{0}+x_{0}+\varepsilon-\displaystyle\frac{y}{2\gamma\tilde{\alpha}^{2}}e^{\int_{0}^{T}-2r_{s}+\left\lVert\kappa_{s}\right\rVert^{2}\mathrm{d}s}\left(\Phi\left(d_{2}\left(\xi_{3}^{*},0\right)\right)-\Phi(d_{2}(\alpha\hat{\xi},0))+\Phi\left(d_{2}\left(\xi_{1}^{*},0\right)\right)\right),$
    where we used that $k_{2}e^{-\int_{0}^{T}r_{s}\mathrm{d}s}\leq x_{0}$, $\tilde{\alpha}+\alpha_{2}=\alpha$, and $\frac{\tilde{\alpha}}{\alpha}\leq 1$ in the first inequality. Moreover, we used (A.17) and $\Phi\left(d_{1}\left(\xi_{3}^{*},0\right)\right)\leq 1$ for all $y>0$ in the second inequality. Due to $\xi_{3}^{*}>0$ and (A.17), it holds that $\Phi\left(d_{2}\left(\xi_{3}^{*},0\right)\right)-\Phi(d_{2}(\alpha\hat{\xi},0))+\Phi\left(d_{2}\left(\xi_{1}^{*},0\right)\right)>0$. Hence, the claim follows, i.e., $\lim_{\varepsilon\to 0}y_{\lambda_{\varepsilon}}^{*}=0$ by the same argument as in Case 1. Note that the existence of a solution was already ensured in the main part of the proof. ∎
@@
    We prove this by contradiction. Therefore, we assume that there exists an $L>0$ such that $\limsup_{\lambda\to\infty}y^{*}_{\lambda}\leq L$. Under this assumption, it holds that $\xi_{1}^{*}\xrightarrow{\lambda\to\infty}\infty$ since $y_{\lambda}^{*}$ is bounded and $\lambda-\sqrt{l(\lambda)}\xrightarrow{\lambda\to\infty}\infty$ for all affine functions $l$. Then (A.17) implies that $\hat{\xi},\bar{\xi},\xi_{2}^{*},\xi_{3}^{*}\xrightarrow{\lambda\to\infty}\infty$ and we know that $\Phi(d_{1}(\xi_{3}^{*},0))\geq\Phi(d_{1}(\alpha\hat{\xi},0))$. However, then $\lim_{\lambda\to\infty}\inf_{0\leq y\leq L}f_{1}(y,\lambda)=\infty$ which is a contradiction to $y^{*}_{\lambda}$ being the zero root of $f_{1}$. ∎